OpenStudy (anonymous):

Given f ′(x) = (5 – x)(8 – x), determine the intervals on which f(x) is increasing or decreasing. a)Decreasing (–∞, 5); increasing on (8, ∞) b)Increasing (–∞, –5) U (–8, ∞); increasing on (–5, –8) c)Decreasing (5, 8); increasing on (–∞, 5) U (8, ∞) d)Decreasing (–∞, 5) U (8, ∞); increasing on (5, 8)

OpenStudy (michele_laino):

Please try to solve the inequality below: \[\left( {5 - x} \right)\left( {8 - x} \right) \geqslant 0\]

OpenStudy (anonymous):

do you want me to factor it out to get: 40-5x-8x+x^2 40-13x+x^2

OpenStudy (anonymous):

@Nnesha @ganeshie8 @iambatman @EclipsedStar

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Kainui

OpenStudy (michele_laino):

for example, try to solve this: 5-x >=0

OpenStudy (anonymous):

oh so x= -5 and x = -8

OpenStudy (michele_laino):

no, those are the roots, you have to find the solution of the inequality (5-x)(8-x)>=0, instead

OpenStudy (anonymous):

i dont see where the > comes in place with the = sign

OpenStudy (anonymous):

here i'll give it a shot: 5-x>=0 -x>=-5 x>=5

OpenStudy (michele_laino):

there is a theorem which states this: " a function f(x) is an increasing function if and only if f '(x) >=0" So you have to apply that theorem

OpenStudy (anonymous):

do you mean greater than or equal to 0?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

the last step is incorrect since if you multiply an inequality by a negative quantity, then that inequality change its direction, so if we have -x >= -5 then x <= 5

OpenStudy (anonymous):

ohhh i thought you had them as 2 separate things, perfect :) yes im sorry i forgot that rule so i dont see how it relates to the answer

OpenStudy (michele_laino):

it is simple, please solve this another inequality: 8-x >=0

OpenStudy (anonymous):

x<=8

OpenStudy (anonymous):

i think i figured it out! is it C? :)

OpenStudy (michele_laino):

that's right so, we have this subsequent drawing:

OpenStudy (michele_laino):

wait please!

OpenStudy (anonymous):

c)Decreasing (5, 8); increasing on (–∞, 5) U (8, ∞)

OpenStudy (michele_laino):



OpenStudy (michele_laino):

do you unerstand that diagram?

OpenStudy (anonymous):

yes it is increasing

OpenStudy (michele_laino):

oops...understand...

OpenStudy (anonymous):

on (5,8)

OpenStudy (michele_laino):

that diagram tell us that when x<=5 both quantities namely (5-x) and (8-x) are positive, so their product is also positive

OpenStudy (anonymous):

got it

OpenStudy (michele_laino):

now please look at this diagram:

OpenStudy (michele_laino):



OpenStudy (michele_laino):

that diagram tell usthat whe x is grether than 8 both quantities, namely (5-x) and (8-x) are neative, nevertheless their product is stil positive.

OpenStudy (michele_laino):

oops ...are negative...

OpenStudy (anonymous):

its okay :) with that being said, i think its b because positive means increase correct?

OpenStudy (michele_laino):

please wait our conclusion is that the function f'(x) is positive when x <5 and x >8 so applying the theorem above, our function f(x) is an icreasing function when x<5 and x>8

OpenStudy (michele_laino):

yes! first derivative positive, means increasig function

OpenStudy (anonymous):

okay so is it b? :) b)Increasing (–∞, –5) U (–8, ∞); increasing on (–5, –8)

OpenStudy (michele_laino):

I think that ur function is an incresing function in \[\left( { - \infty ,5} \right) \cup \left( {8, + \infty } \right)\]

OpenStudy (anonymous):

ohh true cause it would be the opposite decreasing on both positive 5 & 8 (5,8) :) THANKU!!

OpenStudy (michele_laino):

thank you!!

OpenStudy (anonymous):

U r an amazing teacher!!!! :D

OpenStudy (michele_laino):

thank! again!